# SkillWrapper: Generative Predicate Invention for Skill Abstraction

### v3

@@ B.2 Predicate Invention
4:  while $\exists\langle s_{i},\underline{\omega}_{i},s^{\prime}_{i}\rangle,\langle s_{j},\underline{\omega}_{j},s^{\prime}_{j}\rangle\in{\mathcal{D}}\text{ s.t. }s_{i}\in\bm{\alpha}_{\underline{\omega}_{i}},s_{j}\in\bm{\alpha}_{\underline{\omega}_{j}}\text{, but }s_{i}\in{\mathcal{I}}_{\underline{\omega}_{i}},s_{j}\notin{\mathcal{I}}_{\underline{\omega}_{j}}$ do
@@
7:  end while
8:  while $\exists\langle s_{i},\underline{\omega}_{i},s^{\prime}_{i}\rangle,\langle s_{j},\underline{\omega}_{j},s^{\prime}_{j}\rangle\in{\mathcal{D}}\text{ s.t. }s_{i}^{\prime}\in\bm{\zeta}_{\underline{\omega}_{i}},s_{j}^{\prime}\in\bm{\zeta}_{\underline{\omega}_{j}},\text{ but }s_{i}\in{\mathcal{I}}_{\underline{\omega}_{i}},s_{j}\notin{\mathcal{I}}_{\underline{\omega}_{j}}$ do
9:   $\sigma\leftarrow\textsc{NewPredicate}$
10:   $\mathcal{P}\leftarrow\mathcal{P}\,\|\,\sigma$ if ScoreEff($\sigma,\mathcal{P},\omega,\mathcal{D}$)
11:  end while
12:end for
13:return $\mathcal{P}$
Algorithm 6 Scoring Functions for Invented Predicates
1:Input: New predicate $\sigma$, existing predicate set $\mathcal{P}$, skill $\omega$, and skill execution traces $\mathcal{D}$.
2:Parameters: Threshold $h$
3:
4:ScorePrecond:
[…]
9:for $\langle\mathrm{s},\underline{\omega},\mathrm{s}^{\prime}\rangle\ \in\mathcal{D}$ do
10:  if $\exists\,a^{\prime}\in\mathcal{A}^{\prime},\mathbf{o}\subseteq\mathcal{O}\text{ , s.t. }\mathrm{s}\in\textsc{Pre}_{\underline{a}^{\prime}}\text{ , and }\mathrm{s}\in I_{\underline{\omega}}$ then
11:   $\mathrm{valid=valid}+1$
12:  end if
13:  $\mathrm{total=total}+1$
14:end for
15:return $\mathrm{valid}/\mathrm{total}>h$
16:
17:ScoreEff:
[…]
22:for $\langle\mathrm{s},\underline{\omega},\mathrm{s}^{\prime}\rangle\ \in\mathcal{D}$ do
23:  if $\exists\,a^{\prime}\in\mathcal{A^{\prime}},\mathbf{o}\subseteq\mathcal{O}\text{ , s.t. }\Gamma_{\mathcal{P^{\prime}}}(\mathrm{s^{\prime}})\setminus\Gamma_{\mathcal{P^{\prime}}}(\mathrm{s})=\textsc{Eff}_{\underline{a}^{\prime}}\text{ , and }\mathrm{s}\in I_{\underline{\omega}}$ then
24:   $\mathrm{valid=valid}+1$
25:  end if
26:  $\mathrm{total=total}+1$
27:end for
28:return $\mathrm{valid}/\mathrm{total}>h$
B.3 Operator Learning
Algorithm 7 Learn Operators
1:Input: Skill execution traces $\mathcal{D}(\omega)=\{\langle\mathrm{s},\underline{\omega},\mathrm{s}^{\prime}\rangle\}_{\omega}$, predicate set $\mathcal{P}$
2:Output: Operator set $\mathcal{A}$
[…]
4:for $\langle\mathrm{s},\underline{\omega},\mathrm{s}^{\prime}\rangle\ \in\mathcal{D}$ do
5:  $\mathrm{eff}\leftarrow\Gamma(\mathrm{s}^{\prime})\setminus\Gamma(\mathrm{s})$
6:  $\mathrm{eff\_dict}[\mathrm{eff}]\leftarrow\mathrm{eff\_dict}[\mathrm{eff}]\,\|\,(\,\mathrm{s},\mathrm{s}^{\prime})$
7:end for
[…]
9:for $\mathrm{eff}\in\mathrm{eff\_dict}$ do
10:  $\mathrm{execution\_list}\leftarrow\mathrm{eff\_dict}[\mathrm{eff}]$
11:  $\mathrm{precond}\leftarrow\Pi_{\langle\mathrm{s},\underline{\omega},\mathrm{s}^{\prime}\rangle\ \in\mathrm{execution\_list}}\Gamma(\mathrm{s})$
12:  $\mathcal{A}\leftarrow\mathcal{A}\,\|\,[\mathrm{precond},\mathrm{eff}]$
13:end for
14:return $\mathcal{A}$
Appendix C Properties of Learned Symbolic Models
Relational predicates are the basic units of the abstract representation of the low-level state space. In this section, we characterize the conditions of the learned representations of a finite set of skills using relational predicates to support high-level planning, in the context of model learning. From here, chaining the skills is enabled by applying predicates from the representation of each low-level skill to others for clustering (described in Section 3.3).
Effective skill planning requires an accurate abstract model and grounding function. All forms of abstractions are typically lossy, i.e., while learning an abstract transition model, certain low-level environment details may not be captured. Conversely, the learned model must accurately retain the information needed to produce sound and complete plans. In this section, we characterize the conditions under which an abstract model facilitates exact, sound, and/or complete planning, for the purpose of constraining how we can construct such a model.
@@ Appendix C Properties of Learned Symbolic Models
To begin formalizing the relationship between a skill $\omega\in\Omega$ and some abstract model $\mathcal{M}=({\mathcal{P},\mathcal{A}})$, we define two sets, $\bm{\alpha}_{\omega}$ and $\bm{\zeta}_{\omega}$, representing the states in which the model predicts that the skill may either be initiated (when $s\in\bm{\alpha}_{\omega}$) or terminated (when $s\in\bm{\zeta}_{\omega}$), respectively.
Definition 3.
Given a skill instance $\underline{\omega}\in\underline{\Omega}$ and abstract model $\mathcal{M}=({\mathcal{P},\mathcal{A}})$, we define $\bm{\alpha}_{\underline{\omega}}$ and $\bm{\zeta}_{\underline{\omega}}$ as follows:
$\bm{\alpha}_{\underline{\omega}}$ $=\bigcup_{\underline{a}\in\underline{\mathcal{A}}_{\underline{\omega}}}\mathcal{G}({\underline{\textsc{Pre}}_{\underline{a}}})$ (2)
[…]
For a non-instantiated skill $\omega\in\Omega$, we define $\bm{\alpha}_{\omega}=\cup_{\underline{\omega}}\bm{\alpha}_{\underline{\omega}}$ and $\bm{\zeta}_{\omega}=\cup_{\underline{\omega}}\bm{\zeta}_{\underline{\omega}}$.
We define an exact abstract model as one that perfectly captures the initiation set and termination set of all skills. Although such a representation is infeasible to learn in practice, its properties provide an “ideal case” from which other definitions can weaken assumptions.
Definition 4 (Exact Model).
Let ${\mathcal{M}}=\left({\mathcal{P}},{\mathcal{A}}\right)$ be a model for environment $({\mathcal{S},\mathcal{T},\Omega,T})$ and the set of skills $\Omega$, and let $\alpha_{\omega}$ and $\zeta_{\omega}$ be the approximate initiation and termination sets (Def. 3). The model $\mathcal{M}$} is an exact model iff:
$\forall\omega\in\Omega,s\in\mathcal{S}\colon s\in\mathcal{I}_{\omega}\iff s\in\bm{\alpha}_{\omega}\text{ and }\beta_{\omega}(s)=1\iff s\in\bm{\zeta}_{\omega}.$ (4)
An exact planning model supports accurate planning as it precisely characterizes skills’ initiation and termination sets. However, this accuracy comes at the cost of practical feasibility, as any exact model achieves very little in terms of abstraction: it must express the full initiation and termination sets of each skill. Therefore, in many settings, alternative model properties that approximate the exactness of the learned model, namely soundness, suitability, and correctness, may be preferred as objectives for model learning. We now define these properties.
A symbolic model is sound if it correctly predicts the effects of a plan: whenever a complete and sound planner predicts that a sequence of skills will reach some abstract state, executing the corresponding skills in the environment truly leads there. Soundness rules out spurious symbolic transitions that do not correspond to realizable outcomes. Formally, we define soundness as follows: Let $\mathcal{M}=(\mathcal{P},\mathcal{A})$ denote a symbolic planning model, where $\mathcal{P}$ is a finite set of predicates defining an abstract state space $\bar{\mathcal{S}}$, and $\mathcal{A}$ is a set of abstract actions (skills) with preconditions and effects expressed in terms of $\mathcal{P}$. Each abstract state $\bar{s}\in\bar{\mathcal{S}}$ is obtained by a learned grounding function $\Gamma:\mathcal{S}\to\bar{\mathcal{S}}$ that maps low-level agent states $s\in\mathcal{S}$ to truth assignments over $\mathcal{P}$.
Definition 5 (Soundness).
The model $\mathcal{M}$ is sound iff, for any valid plan $\pi$ produced by a complete symbolic planner over $\mathcal{M}$ and for all task instances $\mathbf{p}_{i}\in\mathbf{P}$,
$\Gamma\!\left(\mathcal{T}(\pi,s_{0})\right)\;=\;\bar{\mathcal{T}}_{\mathcal{M}}\!\left(\pi,\Gamma(s_{0})\right),$
where $s_{0}$ is the initial state, $\mathcal{T}(\pi,s_{0})$ is the set of states reachable by executing $\pi$ from $s_{0}$, and $\bar{\mathcal{T}}_{\mathcal{M}}(\pi,\Gamma(s_{0}))$ is the abstract state predicted by $\mathcal{M}$ after executing $\pi$ from $\Gamma(s_{0})$.
A symbolic model is complete if it never omits real solutions: whenever the environment admits a way to solve a task, the planner can find a corresponding abstract plan in the model. Completeness rules out gaps in symbolic coverage that would make feasible problems appear unsolvable.
Definition 6 (Completeness).
The model $\mathcal{M}=(\mathcal{P},\mathcal{A})$ is complete if, for any task instance $\mathbf{p}_{i}\in\mathbf{P}$ and any sequence of low-level actions that achieves the goal from an initial state $s_{0}$, there exists a symbolic plan $\pi$ over $\mathcal{M}$ such that
$\Gamma\!\left(\mathcal{T}(\pi,s_{0})\right)\models G_{i},$
where $G_{i}$ is the goal condition of $\mathbf{p}_{i}$.
A symbolic model is suitable if it correctly characterizes when a skill can be applied, meaning the symbolic preconditions predicted by the model align with the skill’s real initiation conditions. A skill is applicable in an abstract state iff it is applicable in the corresponding grounded state.
Definition 7 (Suitability).
The model $\mathcal{M}$ is suitable if, for any valid plan $\pi$ produced by a complete symbolic planner and for all task instances $\mathbf{p}_{i}\in\mathbf{P}$,
$\Gamma\!\left(\mathcal{T}(\underline{a_{i}},s_{0})\right)\in\bar{I}_{\underline{a_{i+1}}}\;\;\Longleftrightarrow\;\;\mathcal{T}(\omega_{i},s_{0})\in I_{\omega_{i+1}},\quad\forall a\in\mathcal{A},\forall\omega\in\Omega$
where $s_{0}$ is an initial agent state, $\mathcal{T}(\pi,s_{0})$ is the state reached by executing $\pi$ from $s_{0}$, and $\bar{I}_{a}$ is the abstract initiation set of abstract skill $a$.
The minimum requirement for a model $\mathcal{M}$ to solve abstract planning problems is that it is suitable and complete, such that an abstract plan can always be found, and that it is always executable, if there exists a skill plan as a solution. Although methods for constructing such models exist and have been investigated in previous work, they do not provide these guarantees.
Appendix D Proofs
Lemma 2 (Predicate invention).
Let $\mathcal{D}$} be the set of transition tuples $\langle s,\omega,s^{\prime}\rangle$ with $s,s^{\prime}\in{\mathcal{S}}$ and $\omega\in\Omega$. Let $B_{n}$ be a buffer of $n$ i.i.d. samples from $\mathcal{D}$}, each labeled with outcome and containing at least one successful and one failed transition. Let ${\mathcal{M}}_{n}=({\mathcal{P}}_{n},{\mathcal{A}}_{n})$ be the model learned by SkillWrapper from $B_{n}$, where each operator in ${\mathcal{A}}_{n}$ corresponds to some skill $\omega\in\Omega$. Then, for every $\omega\in\Omega$ appearing in $B_{n}$,
$I_{\omega}\ \triangle\ \bm{\alpha}_{\omega}$ $=\emptyset,$ (5)
$\beta_{\omega}\ \triangle\ \bm{\zeta}_{\omega}$ $=\emptyset,$ (6)
[…]
Proof sketch.
Suppose for contradiction that $\exists\langle s,\omega,s^{\prime}\rangle\in B_{n}$ with $s\in I_{\omega}\triangle\bm{\alpha}_{\omega}$. Two cases arise:
(i) False positive. $s\notin I_{\omega}$ but $s\in\mathbf{\alpha}_{\omega}$. By construction of $\alpha_{\omega}$, there must also exist $s_{j}\in I_{\omega}\cap\alpha_{\omega}$. Thus $1_{\alpha_{\omega}}(s)=1_{\alpha_{\omega}}(s_{j})$ while $I_{\omega}(s)\neq I_{\omega}(s_{j})$, contradicting the update rule.
(ii) False negative. $s\in I_{\omega}$ but $s\notin\alpha_{\omega}$. Then some $s_{j}\notin I_{\omega}\cap\alpha_{\omega}$ must also exist, yielding the same contradiction.
Hence, no such $s$ exists, and Eq. (6) holds. The proof for Eq. (7) is identical, replacing $\alpha_{\omega}$ with $\zeta_{\omega}$ defined from operator effects. ∎

### v4

@@ III-B Predicate Invention
4:  while precondition inconsistency exists do $\triangleright$ equation 1
@@
7:  while effect inconsistency exists do $\triangleright$ equation 2
8:    $\sigma\leftarrow\textsc{NewPredicate}$
9:    $\mathcal{P}\leftarrow\mathcal{P}\,\|\,\sigma$ if ScoreEff($\sigma,\mathcal{P},\omega,\mathcal{D}$)
10:return $\mathcal{P}$
@@
To formally describe the conditions, we define two sets, $\bm{\alpha}_{\omega}$ and $\bm{\zeta}_{\omega}$, representing the states in which the model predicts that the skill may either be initiated (when $s\in\bm{\alpha}_{\omega}$) or terminated (when $s\in\bm{\zeta}_{\omega}$), respectively. Both sets are derived from and defined with the learned operators, and their formal definitions can be found in Appendix -B.
$\begin{split}\exists\langle s_{i},\underline{\omega}_{i},s^{\prime}_{i}\rangle,\langle s_{j},\underline{\omega}_{j},s^{\prime}_{j}\rangle\in\mathcal{D}\text{ s.t. }s_{i}\in\bm{\alpha}_{\underline{\omega}_{i}},s_{j}\in\bm{\alpha}_{\underline{\omega}_{j}},\\ \text{but }s_{i}\in\mathcal{I}_{\underline{\omega}_{i}}\text{ and }s_{j}\notin\mathcal{I}_{\underline{\omega}_{j}}.\end{split}$ (1)
Figure 2: Example of Predicate Invention. The initial states of two transitions are both said to satisfy the preconditions of certain operators learned from the same skill, while transition 1 is successful, but transition 2 is not. In this case, the first condition (equation 1) is triggered, and the foundation model is prompted with both transitions to invent a new predicate.
$\begin{split}\exists\langle s_{i},\underline{\omega}_{i},s^{\prime}_{i}\rangle,\langle s_{j},\underline{\omega}_{j},s^{\prime}_{j}\rangle\in{\mathcal{D}}\text{ s.t. }s_{i}^{\prime}\in\bm{\zeta}_{\underline{\omega}_{i}},s_{j}^{\prime}\in\bm{\zeta}_{\underline{\omega}_{j}},\\ \text{ but }s_{i}\in{\mathcal{I}}_{\underline{\omega}_{i}}\text{ and }s_{j}\notin{\mathcal{I}}_{\underline{\omega}_{j}}.\end{split}$ (2)
Contrastive predicate proposal
Empirical predicate selection
Although foundation models provide a strong prior on which information is skill-relevant for predicate proposal, they may produce errors and hallucinations. To ensure robustness, we introduce a scoring function that estimates the usefulness of a candidate predicate by trying adding it to the vocabulary and learning hypothetical operators. For each successful transition in $\mathcal{D}$}, there must exist at least one operator with preconditions satisfied by the transition’s initial state; for each failed transition, no such operator may exist. Effect evaluation follows the same principle. After evaluating over $\mathcal{D}$, we decide whether to add the candidate predicate based on an empirical threshold. Further details of the scoring function can be found in Algorithm 6 (App. -A).
III-C Operator Learning
Our operator learning procedure extends the associative model learning paradigm [3] to the setting of skill abstraction.
Associative model learning
A single skill may induce multiple distinct abstract state changes depending on the context of execution. To represent these conditional effects, SkillWrapper clusters observed transitions based on their lifted (object-agnostic) effect sets, enabling it to learn a single operator across distinct instantiations of a skill. The preconditions of each operator are then computed as the intersection of all initial abstract states in the corresponding transitions, ensuring that each operator is both minimal and consistent. The learning process operates directly on the abstract level, allowing it to capture relationships over objects beyond the skill parameters. We provide a detailed explanation in App. -G.
Multi-type object-centrism
Predicate re-scoring
III-D Theoretical Analysis
We now provide theoretical guarantees for SkillWrapper. These results show that the learned symbolic model is sound with respect to observed data and converges to a complete model with high probability. Full proofs are deferred to App. -C.
Theorem 1 (Soundness of SkillWrapper).
Lemma 1.